a shelf is built into the corner of a room whose walls meet at an angle of 120 degree. Find the area of the shelf space if its front edge is 24cm and the other two edges are equal

Answer by rothauserc(4718)

we have a triangle with one 120 degree angle and two 30 degree angles. A line which bisects the 120 degree angle bisects the 24 cm edge and is perpendicular to the 24 cm edge, then
tan(30) = h / 12 where h is the height of the triangle
h = tan(30) * 12 = 6.92820323
area of shelf space = 2 * (1/2) * 12 * 6.92820323
area of shelf space = 83.138438763 square cm's